# Exploring Holomorphic Retracts

### v6

@@ Conclusion
Before we begin to work out the proof, we need to make some general remarks; to indicate as to why there are enough instances of domains verifying the conditions stipulated in proposition 1.6. Indeed, even before this, let us indicate as to why it may perhaps be of mild interest atleast, to note in passing (as we did in the preliminaries section) that: one may deduce out of the general facts about retracts and the already mentioned well-known characterization of retracts of polydiscs, yet another proof of the dramatic failure of the Riemann mapping theorem in multidimensional complex analysis which is, the biholomorphic inequivalence of two of the simplest bounded simply connected domains namely, the ball $\mathbb{B}$ and $\Delta^{N}$ (despite their diffeomeorphic equivalence), as soon as the dimension $N$ gets any bigger than one; and other such inequivalences using retracts. We have already put this proof for a record in the ‘preliminaries’ section wherein we have formulated it more generally for a ball $D$ with respect to an arbitrary norm with $\mathbb{C}$-extremal boundary and product domains much more general than polydiscs (and so that proof is not repeated here). However here, we would indeed like to hasten to mention that this does not obtain for such $D$, results as good as those, which already well-known techniques in this realm can! – so, this leads to the question as to why one may want to even consider this proof even if the problem of biholomorphic inequivalence and its special case of establishing irreducibility of a domain $D$ i.e., showing that $D$ is not biholomorphic to a product of lower dimensional domains, is undoubtedly fundamental. To address this briefly here, let us recall that while the well-known techniques alluded to above, more explicitly such as that of Remmert – Stein (exposited well in [Encyclo-Vol6]; more specifically, see theorem $4.1$ therein alongwith the subsequent theorem $4.2$ due to Rischel) for showing irreducibility of domains, work for far more general domains $D$ beyond balanced domains of holomorphy, they impose the condition that $D$ not have any non-trivial analytic varieties in the boundary. We shall presently show that the route via retracts can very well dispense with such conditions, if $D$ is a balanced domain of holomorphy. To exposit this better by comparing and contrasting with the Remmert – Stein techniques, let us mention that the aforementioned proof by retracts for $\mathbb{B}\not\simeq\Delta^{N}$, shares (unlike a proof based on comparison of automorphism groups for instance), with the best possible techniques in this context, the essential idea about the difference in the lack or presence of analytic discs in the boundary of $D$ as being responsible for such inequivalences, as it has a bearing on the geometry of retracts through the origin in $D$. Moreover, this provides a shorter way to establish inequivalences without having to worry about the boundary extendibility (or boundary behaviour) of biholomorphic mappings since the boundary geometry is ‘internalized’ as features about the retracts through points in the interior; we believe this will be useful for establishing new inequivalences between pairs of bounded balanced domains in both of whose boundaries there reside non-trivial analytic discs. To illustrate this point with a concrete and simple example, consider the question of whether the analytic polyhedron
$D_{h}=\{(z,w)\in\mathbb{C}^{2}:|w^{2}-z^{2}|<1\}\cap\{(z,w)\in\mathbb{C}^{2}:|zw|<2\}$
@@
is biholomorphic to a product domain  [^5] A motivation for the specific choices here, is that the simplest examples of bounded product domains namely polydiscs are defined as sublevel sets of a pair of (homogeneous) holomorphic linear polynomials and we are interested in bounded domains whose definition bears some semblance to that of polydiscs, but are nevertheless irreducible.. To begin with analysing this example first, we note that while an open set cut-out by such polynomial inequalities in general may fail to be connected, such issues do not arise here: the homogeneity of the defining pair of quadratic polynomials leads to $D_{h}$ being balanced, thereby (immediately seen to be connected and) contractible. Let us note in passing that $D_{h}$ is not Reinhardt; more importantly, that $D_{h}$ is indeed bounded: to verify this, note that the defining pair of inequalities together leads to
@@
when squared and combined with 8.1, leads to
@@
However, we would like to show this and other modified versions now by circumventing use of the Riemann mapping theorem which may not work in dealing with more complex cases. Indeed, to spoil the above proof which is based on the convenient availability of a model domain in the equivalence class of balanced domains which are biholomorphic to products domains as nice as the bidisc, let us remove an analytic hypersurface $A$ not passing through the origin [^6] Not removing the origin thus, helps in rendering a short illustration of our use of retracts, owing to scarcity of results that we can readily cite and use. It may be noted to begin with here that this modified domain only has more analytic varieties in the boundary, that is to say: the modification to spoil the foregoing proof, has not been done at the cost of the very purpose of the previous example!, and consider $D_{h}^{A}=D_{h}\setminus\{A\}$. While the removal of $A$ from $D_{h}$ never disconnects $D_{h}$, it always destroys the simply connectedness of $D_{h}$, as is always the case when proper analytic sets are subtracted from any domain. Though not required by the statement of proposition 1.6, the case $D=D_{h}$ also serves as a concrete example of a (hyperbolic) balanced domain of holomorphy with analytic discs in its boundary (without purposefully removing analytic sets!), which feature as already mentioned is the case of greatest interest in contributing to any new results or routes about establishing non-existence of a product structure in a domain.
To establish the irreducibility of such $D_{h}^{A}$, let us start with the most natural attempt of assuming to obtain a contradiction, the existence of a biholomorphism $F$ from $D_{h}^{A}$ onto a product domain $D$; no less standard then, is to analyze the boundary behaviour of $F$. However, this is rather cumbersome owing to potential complexity of the boundary of the factors $B_{1},B_{2}$. It is here where consideration of retracts provides a far more convenient approach, as already alluded to in the above.
However, before we discuss the proof of propositions about the irreducibility of analytic complements as in proposition 1.6, it is convenient to formulate part of its proof as the following more general lemma.
@@
Let $D$ be any bounded pseudoconvex domain in $\mathbb{C}^{N}$ either with a Stein neighbourhood basis or with a bounded plurisubharmonic defining function. Then any holomorphic retraction map on $D\setminus A$, where $A$ is a proper analytic subset of $D$ can be extended to a holomorphic retraction map on $D$. Consequently, every retract $R^{A}$ of $D\setminus A$ is of the form $R^{A}=R\cap(D\setminus A)$ for some retract $R$ of $D$.
@@
Firstly recall that removing an analytic set from a domain does not disconnect it; by definition $A$ is closed in $D$, thereby $D\setminus A$ is a domain in $\mathbb{C}^{N}$. Let $\rho:D\setminus A\to D\setminus A$ be a holomorphic retraction map. By applying the Riemann removable singularities theorem (for holomorphic functions of several complex variables) to each of the component functions of $\rho$ we get a (uniquely determined) holomorphic extension of $\rho$ which we denote by $\tilde{\rho}$. The fact that $\rho$ maps $D\setminus A$ into itself implies firstly that $\tilde{\rho}$ maps $D$ into its closure $\overline{D}$. However, it is easy to see that $\tilde{\rho}$ actually maps $D$ into itself. Indeed, pick any $a\in A$ and enquire if it is possible for $b=\tilde{\rho}(a)$ to lie in $\partial D$. Recalling the local structure of analytic sets, we know that: there exist a complex line segment $\ell\subset D$ through $a$ which intersects $A$ only at $a$. If $\varphi:\Delta\to D$ is a parametrization of $\ell$ with $\varphi(0)=a$, then $\tilde{\rho}\circ\varphi$ gives an analytic disc (with its image) intersecting $\partial D$ only at $a$. If $D$ admits a bounded plurisubharmonic defining function $\psi$, then we consider the composition $\phi:=\psi\circ\tilde{\rho}\circ\varphi$, which is a subharmonic function on $\Delta^{*}=\Delta\setminus\{0\}$; as it is also bounded (by the boundedness of $\psi$), it follows by that the origin is a removable singularity for this subharmonic function, thereby $\phi$ extends to a subharmonic function on $\Delta$ which we continue to denote by $\phi$. But then, the maximum principle for this subharmonic function is violated, under our assumption that $b\in\partial D$. We deduce therefore that $b$ cannot lie in $\partial D$ and it follows that $\tilde{\rho}(D)\subset D$. Even if the existence of such a $\psi$ is not known, it is indeed true that $\tilde{\rho}(D)$ lies within $D$ (the above argument argument was given to show a simple route under the aforementioned assumption about bounded plurisubharmonic defining functions). This follows from the last result in [Abe], owing to our hypothesis that $D$ has a Stein neighbourhood basis; as the work [Abe] has some non-major errors, we also cite the Masters thesis of Persson [Persson], wherein the errors are set right with a better exposition. Thus as $\tilde{\rho}(a)$ lies within $D$ and as $a$ was arbitrary, we conclude that $\tilde{\rho}$ maps $A$ (and thereby $D$ as well) into $D$.
@@
One may want to know to what extent we may weaken the hypotheses of the above lemma. We do not pursue this here but draw attention to the fact that rather simple counterexamples exist to show that boundedness cannot be dropped easily. Indeed, consider the unbounded domain $D=\mathbb{C}^{2}\setminus A$, where $A=\{z=0\}$, the $w$-axis. Notice then that $R^{A}=\{(z,1/z):z\in\mathbb{C}^{*}\}$ is a retract of $D$. We observe that there exists no retract $R$ of $\mathbb{C}^{2}$ such that $R^{A}=R\cap(D\setminus A)$. Indeed, this follows from the fact that the (smooth) analytic variety $R^{A}$ described by the equation $zw-1=0$, is irreducible, together with the principle of analytic continuation for complex analytic sets.
@@
Proof of proposition 1.5 . Let $p=(p_{1},p_{2})\in D_{1}\times D_{2}$ Since $D_{1}$ is bounded, there exist $r_{1}>0$ such that $D_{1}\subset\Delta(p_{1},r_{1})$. Since $D_{2}$ is open, there exist $r_{2}>0$ such that $\Delta(p_{2},r_{2})\subset D_{2}$. Consider the domain $D:=D_{1}\times D_{2}$. Choose $v_{2}\in T_{p_{2}}D_{2}$ with $|v_{2}|<\frac{r_{2}}{r_{1}}$. Consider a holomorphic map $F:D_{1}\to\mathbb{C}$ defined by $F(z)=p_{2}+(z-p_{1})v_{2}$. Note that $F(p_{1})=p_{2}$ and $F^{\prime}(p_{1})=v_{2}$. Take $\epsilon=\frac{r_{2}}{r_{1}}$. If $|v_{2}|<\epsilon$, then $F$ maps $D_{1}$ into $D_{2}$. Consider a holomorphic retraction map $G:D_{1}\times D_{2}\to D_{1}\times D_{2}$ defined by $G(z,w)=(z,F(z))$. Hence $Z:=\{(z,F(z)):z\in D_{1}\}$ is a one-dimensional retract of $D_{1}\times D_{2}$ passing through $(p_{1},p_{2})$ and $(1,v_{2})\in T_{p}Z$. ∎
@@
Suppose $D_{1},D_{2}$ be domains in $\mathbb{C}^{N},\mathbb{C}^{M}$ respectively and atleast one of domains is Caratheodory hyperbolic, say $D_{1}$; let $p=(p_{1},p_{2})$ be any given point in the product domain $D=D_{1}\times D_{2}$. Let $v_{1}\in T_{p_{1}}D_{1}$ be any non-zero vector. Then there exist $\epsilon>0$, such that for all vectors $v_{2}\in T_{p_{2}}D_{2}\cong\mathbb{C}^{M}$ with $\|v_{2}\|_{\infty}<\epsilon$, there is a retract $Z$ of $D_{1}\times D_{2}$ passing through $p$ such that $(w_{1},v_{2})\in T_{p}Z$, for some non-zero vector $w_{1}$ with $w_{1}\in span_{\mathbb{C}}\{v_{1}\}$.

### v7

@@ Conclusion
The above exercise in showing irreducibility may seem to need justification, owing to the existence of well-known methods which yield better results for certain classes of domains even if they are different. To put this matter in perspective we recall that the famous inequivalence between the Euclidean ball $\mathbb{B}$ and the polydisc $\Delta^{N}$ for all $N\geq 2$, can be shown using retracts as well; while we believe that this proof will atleast be of mild interest to atleast note in passing, we have better justifications below. Now, we have already put this proof for a record in the ‘preliminaries’ section wherein we have formulated it more generally for a ball $D$ with respect to an arbitrary norm with $\mathbb{C}$-extremal boundary and product domains much more general than polydiscs (so that proof is not repeated here). Here, we would indeed like to hasten to mention that this does not obtain for such $D$: results as good as can be obtained from already well-known techniques in this realm can! – so, this leads to the question as to why one may want to even consider this proof even if the problem of biholomorphic inequivalence and its special case of establishing irreducibility of a domain $D$ i.e., showing that $D$ is not biholomorphic to a product of lower dimensional domains, is undoubtedly fundamental. To address this briefly here, let us recall that while the well-known techniques alluded to above, more explicitly such as that of Remmert – Stein (exposited well in [Encyclo-Vol6]; more specifically, see theorem $4.1$ therein alongwith the subsequent theorem $4.2$ due to Rischel) for showing irreducibility of domains, work for far more general domains $D$ beyond balanced domains of holomorphy, they impose the condition that $D$ not have any non-trivial analytic varieties in the boundary. We shall presently show that the route via retracts can very well dispense with such conditions, if $D$ is a balanced domain of holomorphy. To exposit this better by comparing with the Remmert – Stein techniques, let us mention that the aforementioned proof by retracts for $\mathbb{B}\not\simeq\Delta^{N}$, shares (unlike a proof based on comparison of automorphism groups for instance), with the best possible techniques in this context, the essential idea about the difference in the lack or presence of analytic discs in the boundary of $D$, as being responsible for such inequivalences; this difference has a bearing on the geometry of retracts through the origin in $D$. Moreover, this provides a shorter way to establish inequivalences without having to worry about the boundary extendibility (or boundary behaviour) of biholomorphic mappings since the boundary geometry is ‘internalized’ (atleast partly) as features about the retracts through points in the interior; for instance, holomorphic extremality of the boundary of a bounded balanced domain of holomorphy entails linearity of the retracts through its center and these linear slices have holomorphically extreme boundaries again. We believe this will be useful for establishing new inequivalences between pairs of bounded balanced domains in both of whose boundaries there reside non-trivial analytic discs. We now illustrate these points with a couple of progressively complex examples and theorems. We start with a concrete and simple example of a special analytic polyhedron; indeed, this is the simplest example of a bounded balanced analytic polyhedron with $\mathbb{C}$-extremal boundary and will serve as a running example for us. Specifically, consider the question of whether
(8.1) $D_{h}:=\{(z,w)\in\mathbb{C}^{2}:|w^{2}-z^{2}|<1\}\cap\{(z,w)\in\mathbb{C}^{2}:|zw|<2\}$
@@
is biholomorphic to a product domain  [^11] A motivation for the specific choices here, is that the simplest examples of bounded product domains namely polydiscs are defined as sublevel sets of a pair of (homogeneous) holomorphic linear polynomials and we are interested in bounded domains whose definition bears some semblance to that of polydiscs, but are nevertheless irreducible. Indeed as will be seen, this example is the simplest special analytic polyhedron which is balanced and non-linear.. To begin with analysing this example first, we note that while an open set cut-out by such polynomial inequalities in general may fail to be connected, such issues do not arise here: the homogeneity of the defining pair of quadratic polynomials leads to $D_{h}$ being balanced, thereby (immediately seen to be connected and) contractible. Let us note in passing that $D_{h}$ is not Reinhardt; more importantly, that $D_{h}$ is indeed bounded: to verify this, note that the defining pair of inequalities together leads to
@@
when squared and combined with 8.2, leads to
@@
However, we would like to show this and other modified versions now by circumventing use of the Riemann mapping theorem which may not work in dealing with more complex cases. Indeed, to spoil the above proof which is based on the convenient availability of a model domain in the equivalence class of balanced domains which are biholomorphic to products domains as nice as the bidisc, let us remove an analytic hypersurface $A$ not passing through the origin [^12] Not removing the origin thus, helps in rendering a short illustration of our use of retracts, owing to scarcity of results that we can readily cite and use. It may be noted to begin with here that this modified domain only has more analytic varieties in the boundary, that is to say: the modification to spoil the foregoing proof, has not been done at the cost of the very purpose of the previous example!, and consider $D_{h}^{A}=D_{h}\setminus\{A\}$. While the removal of $A$ from $D_{h}$ never disconnects $D_{h}$, it always destroys the simply connectedness of $D_{h}$, as is always the case when proper analytic sets are subtracted from any domain.
To establish the irreducibility of such $D_{h}^{A}$, let us start with the most natural attempt of assuming to obtain a contradiction, the existence of a biholomorphism $F$ from $D_{h}^{A}$ onto a product domain $D$; no less standard then, is to analyze the boundary behaviour of $F$. However, this is rather cumbersome owing to potential complexity of the boundary of the factors $B_{1},B_{2}$. It is here where consideration of retracts provides a far more convenient approach, as already alluded to in the above, particularly for theorem 1.5, one of the main theorems of this section.
However, before we discuss the proof of theorem 1.5, it is convenient to formulate part of its proof as the following more general lemma.
@@
Let $D$ be any bounded hyperconvex domain in $\mathbb{C}^{N}$. Let $A$ be any non-trivial analytic subset of $D$. Then any holomorphic retraction map on $D\setminus A$, be extended to a holomorphic retraction map on $D$. Consequently, every retract $R^{A}$ of $D\setminus A$ is of the form $R^{A}=R\cap(D\setminus A)$ for some retract $R$ of $D$.
@@
Firstly, recall that removing an analytic set from a domain does not disconnect it; by definition $A$ is closed in $D$, thereby $D\setminus A$ is a domain in $\mathbb{C}^{N}$. Let $\rho:D\setminus A\to D\setminus A$ be a holomorphic retraction map. By applying the Riemann removable singularities theorem (for holomorphic functions of several complex variables) to each of the component functions of $\rho$ we get a (uniquely determined) holomorphic extension of $\rho$ which we denote by $\tilde{\rho}$. The fact that $\rho$ maps $D\setminus A$ into itself implies firstly that $\tilde{\rho}$ maps $D$ into its closure $\overline{D}$. However, it is easy to see that $\tilde{\rho}$ actually maps $D$ into itself. Indeed, pick any $a\in A$ and enquire if it is possible for $b=\tilde{\rho}(a)$ to lie in $\partial D$. Recalling the local structure of analytic sets, we know that: there exist a complex line segment $\ell\subset D$ through $a$ which intersects $A$ only at $a$. If $\varphi:\Delta\to D$ is a parametrization of $\ell$ with $\varphi(0)=a$, then $\tilde{\rho}\circ\varphi$ gives an analytic disc (with its image) intersecting $\partial D$ only at $a$. Since $D$ is hyperconvex, it follows from the theorem 2.50 and its remark that $\tilde{\rho}(a)$ is a local psh barrier point of $D$. So, there exists neighborhood $U$ of $b:=\tilde{\rho}(a)$ and plurisubharmonic function $\psi$ on $U\cap D$ such that $\psi<0$ on $U\cap D$ and $\lim_{U\cap D\ni z\to b}\psi(z)=0$. After shrinking the radius of the punctured disc $\Delta^{*}$ (if needed), we may assume that $image(\tilde{\rho}\circ\varphi)\subset U$. Consider the subharmonic function $\phi:\Delta^{*}\to\mathbb{R}$ defined by $\phi(z)=\psi\circ\tilde{\rho}\circ\varphi(z)$. By applying the removable singularity theorem for subharmonic functions, we obtain that 0 is a removable singularity for $\phi$. Hence $\phi$ extends to a subharmonic function on $\Delta$ which we continue to denote by $\phi$. By the limiting behaviour of $\psi$ at $b$, we obtain that $\phi(0)=0$. But then, the maximum principle for this subharmonic function is violated, under our assumption that $b\in\partial D$. We deduce therefore that $b$ cannot lie in $\partial D$ and it follows that $\tilde{\rho}(D)\subset D$.
@@
We first recall that for the class of much interest in this article namely bounded balanced pseudoconvex domains, the hypothesis of the above is not satisfied automatically as hyperconvexity of such domains is equivalent to continuity of the Minkowski functional (see Remark 2.45 (iii) in the preliminaries section); nevertheless, it turns out that the conclusion of the above lemma is true regardless of such regularity, owing to certain results that we briefly touch upon now. For our applications however, we shall not require such more general observations. To remark on the extension then, the above lemma can be derived for bounded pseudoconvex domains $D$ even if the existence of such a $\psi$ as in its proof above is not known, provided $D$ has a Stein neighbourhood basis. The only point to be checked is $\tilde{\rho}(D)\subset D$. It is indeed true that $\tilde{\rho}(D)$ lies within $D$ and can be derived rather quickly from known results (the arguments above were given to show a simple route under the aforementioned assumption about bounded plurisubharmonic defining functions). This follows from the last result in [Abe], owing to our hypothesis that $D$ has a Stein neighbourhood basis; as the work [Abe] has some non-major errors, we also cite the Masters thesis of Persson [Persson], wherein the errors are set right with a better exposition. Thus as $\tilde{\rho}(a)$ lies within $D$ and as $a$ was arbitrary, we conclude that $\tilde{\rho}$ maps $A$ (and thereby $D$ as well) into $D$.
Remark 8.3.
One may want to know to what extent we may weaken the hypotheses of the above lemma. We do not pursue this here but draw attention to the fact that rather simple counterexamples exist to show that boundedness cannot be dropped easily. Indeed, consider the unbounded domain $D=\mathbb{C}^{2}\setminus A$, where $A=\{z=0\}$, the $w$-axis. Notice then that $R^{A}=\{(z,1/z):z\in\mathbb{C}^{*}\}$ is a retract of $D$. We observe that there exists no retract $R$ of $\mathbb{C}^{2}$ such that $R^{A}=R\cap D$. Indeed, this follows from the fact that the (smooth) analytic variety $R^{A}$ described by the equation zw1=0zw-1=0, is irreducible, together with the principle of analytic continuation for complex analytic sets.
@@
Proof of proposition 1.4 . Let $p=(p_{1},p_{2})\in D_{1}\times D_{2}$ . Since $D_{1}$ is bounded, there exist $r_{1}>0$ such that $D_{1}\subset\Delta(p_{1},r_{1})$. Since $D_{2}$ is open, there exist $r_{2}>0$ such that $\Delta(p_{2},r_{2})\subset D_{2}$. Consider the domain $D:=D_{1}\times D_{2}$. Choose $v_{2}\in T_{p_{2}}D_{2}$ with $|v_{2}|<\frac{r_{2}}{r_{1}}$. Consider a holomorphic map $F:D_{1}\to\mathbb{C}$ defined by $F(z)=p_{2}+(z-p_{1})v_{2}$. Note that $F(p_{1})=p_{2}$ and $F^{\prime}(p_{1})=v_{2}$. Take $\epsilon=\frac{r_{2}}{r_{1}}$. If $|v_{2}|<\epsilon$, then $F$ maps $D_{1}$ into $D_{2}$. Consider a holomorphic retraction map $G:D_{1}\times D_{2}\to D_{1}\times D_{2}$ defined by $G(z,w)=(z,F(z))$. Hence $Z:=\{(z,F(z)):z\in D_{1}\}$ is a one-dimensional retract of $D_{1}\times D_{2}$ passing through $(p_{1},p_{2})$ and $(1,v_{2})\in T_{p}Z$. ∎
@@
Suppose $D_{1},D_{2}$ be domains in $\mathbb{C}^{N},\mathbb{C}^{M}$ respectively and at least one of domains is Caratheodory hyperbolic, say $D_{1}$; let $p=(p_{1},p_{2})$ be any given point in the product domain $D=D_{1}\times D_{2}$. Let $v_{1}\in T_{p_{1}}D_{1}$ be any non-zero vector. Then there exist $\epsilon>0$, such that for all vectors $v_{2}\in T_{p_{2}}D_{2}\cong\mathbb{C}^{M}$ with $\|v_{2}\|_{\infty}<\epsilon$, there is a retract $Z$ of $D_{1}\times D_{2}$ passing through $p$ with $T_{p}Z\ni(w_{1},v_{2})$, for some non-zero vector $w_{1}$ with $w_{1}\in\operatorname{span}_{\mathbb{C}}\{v_{1}\}$.